### v1

# Classical periodic orbits in extended phase space for spherical harmonic oscillator with spin-orbit coupling

Abstract
A complete analysis of classical periodic orbits (POs) and their bifurcations was conducted in spherical harmonic oscillator system with spin-orbit coupling. The motion of the spin is explicitly considered using the spin canonical variables derived by semiclassical approximation to the spin coherent state path integral representation. In addition to the diametric and two circular PO families with frozen spin, solutions that bridge two circular POs are found in which orbital motion is coupled to spin precession. In addition, each bridge encounters a secondary bifurcation on the way from one circular PO to the other and generates a new PO, that survives at higher energies while maintaining a constant period. The generic expressions for those POs are obtained explicitly, and all the above peculiar bifurcation scenarios are described fully analytically.
I Introduction
Classical periodic orbits (POs) play central roles in semiclassical quantization for both integrable and nonintegral Hamiltonian systems[1, 2, 3, 4]. Gutzwiller derived a famous trace formula that represents the quantum density of states in terms of the contributions of classical POs[5], based on the semiclassical approximation to the path integral representation using the stationary phase method. The formula works paricularly well for strongly chaotic systems. In another approach, Balian and Bloch derived the semiclassical density of states for infinite well potential systems of arbitrary boundary shapes[6]. It is based on the multiple reflection expansion of the transition amplitude, and can be applied to both integrable and nonintegrable models. For integrable Hamiltonian systems whose energy spectra can be approximately obtained by the Einstein-Brillouin-Keller quantization rule (an extension of the Bohr-Sommerfeld quantization condition to multiply-periodic systems of arbitrary dimensions), Berry and Tabor showed that the oscillating part of the level density is generally represented in the form of the periodic orbit sum[7]. The extension of the Gutzwiller formula to the case of continuous symmetries was conducted by Creagh and Littlejohn[8, 9]. Its extension to bifurcation regions, where the original form of the Gutzwiller formula breaks down, was proposed by Sieber et al. using the uniform approximations[10, 11, 12, 13] and by Magner et al. using the improved stationary-phase method [14, 15]. In general, semiclassical approximation of quantum mechanical level density
$g_{\rm qm}(E)=\sum_{n}\delta(E-E_{n})$
for a conservative Hamiltonian system is given by a trace formula expressed in the form
$g_{\rm cl}(E)=\bar{g}(E)+\sum_{\rm PO}A_{\rm PO}\cos\left(\tfrac{1}{\hbar}S_{\rm PO}-\tfrac{\pi}{2}\mu_{\rm PO}\right).$ (1)
[…]
One of the most significant advantages in the trace formula is that it can explain the origin of the regular oscillations in level distribution, which is generally cannot be understood by purely quantum mechanical concepts. It is especially useful for analyzing gross shell structures in finite quantum systems. It has been applied to nuclear deformation soon after its derivation[16, 17], and then to various quantum systems such as metallic clusters[18, 19, 20] and quantum dots[21]. Periodic-orbit theory (POT) has also been applied to the problems of quantum transport. The oscillation of magnetoresistance in mesocopic conductors, which is related to the electron density of states at the Fermi energy, is successfully explained by the contributing POs to the semiclassical level density[22, 23]. The significance of PO bifurcations as the microscopic origin of the deformed shell structure has been emphasized in [20, 24, 14, 25, 26, 27, 28], and its effect on the magnetoresistance was discussed in [21].
In addition to the orbital degrees of freedom, quantum particles have a spin degree of freedom. The Dirac Hamiltonian for a charged fermion in an electromagnetic field includes the spin-orbit potential, whose effect can be measured as the spin-orbit splitting of the atomic energy levels. In atomic nuclei, strong spin-orbit coupling is inevitable in the mean field of a nucleon to account for the nuclear “magic numbers”, numbers of constituent particles (neutrons or protons) of closed-shell configurations[29, 30, 31]. Thus, the inclusion of spin in semiclassical POT is important for a deeper understanding of nuclear shell structures and related phenomena.
To deal with spin-orbit coupling in semiclassical theory, two different prescriptions have been proposed. The first one is the coupled-channel WKB method[32, 33] in which the multicomponent wave equation is diagonalized using the Wigner-Weyl calculus and is reduced to a set of uncoupled equations corresponding to the individual channels. This method was applied to the problem of the spin-orbit coupling[34] by treating the spin as a channel. This is a kind of adiabatic approximation in which spin is regarded as a slow variable. The semiclassical trace formula is then constructed in terms of the POs in those adiabatic Hamiltonians that correspond to individual spin channels. The weakness of this prescription lies in the occurrence of the so-called mode conversion. Adiabatic approximation cannot be justified in the mode conversion region where two adiabatic Hamiltonians cross with each other and classical dynamics become singular. To deal with this problem, a use of diabatic representation has been proposed, which allows a spin flip at the mode conversion point to remove the singularity[35]. It seems working well in some numerical applications, however, it remains ambiguous and lacks theoretical basis[36].
In another prescription, the spin degree of freedom is explicitly incorporated into the equations of motion (EOM). The classical phase space is extended by introducing a spin canonical variables. The classical dynamics of the spin are derived by a semiclassical approximation of the spin coherent state path integral representation[37]. Theoretically, the semiclassical approximation is justified for large spins, but in practice, it also gives reasonable results for small spins in many cases. Thus, we can expect it to provide a better prescription for a semiclassical treatment of the spin, free from the problems like mode conversions. The compensation is the increased degrees of freedom, which usually complicate classical motions. Pletyukov and Zeitsef derived a trace formula which explicitly consider the spin motion[38]. They also showed that the formula works well for small spins (such as $s=\hbar/2$ for electrons and nucleons) in case Hamiltonians depend linearly on the spin. Thus, the proposed method can be safely applied to the spin-orbit problem in electron and nucleon systems. The application to the 2-dimensional Rashba model to electrons in quantum dots was performed in [36, 39].
In this paper, I investigate the 3-dimensional isotropic harmonic oscillator model with a simple spin-orbit coupling term:
$H=\frac{{\bm{p}}^{2}}{2M}+\frac{M\omega^{2}{\bm{r}}^{2}}{2}-\kappa({\bm{r}}\times{\bm{p}})\cdot{\bm{s}}.$ (2)
This Hamiltonian is physically important as the mean field model for light nuclei, corresponding to the Nilsson model without $l^{2}$ term[40, 31]. Classical dynamics in the same Hamiltonian have been studied numerically in [41]. Quantum energy eigenvalue problem for the Hamiltonian (2) is exactly solved, and the energy eigenvalues for particles with spin $s=\hbar/2$ are expressed as
[…]
$(\mathring{\kappa}=\kappa s/\omega,\quad n=0,1,2,\cdots,\quad\mathring{l}=0,1,2,\cdots)$
where $n,\mathring{l},\mathring{\jmath}$ are radial, spin, orbital angular and total angular momentum quantum numbers, respectively. Each level specified by $\mathring{\jmath}$ has $(2\mathring{\jmath}+1)$-fold degeneracies because the energy eigenvalues are independent on the magnetic quantum number $m_{j}$ ($-\mathring{\jmath}\leq m_{j}\leq\mathring{\jmath}$).
For a system in which the energy eigenvalues are explicitly given in terms of the quantum numbers, Berry-Tabor formula[7] can be applied as the semiclassical density of states, and the following formula has been obtained in Ref. [36] (see also Appendix A)
$g(E)=\bar{g}(E)+\sum_{\sigma=\pm}\sum_{L=1}^{\infty}\delta g_{L}^{(\sigma)}(E)+\sum_{N=1}^{\infty}\delta g_{N}^{(0)}(E),$ (6)
$\begin{split}\delta g_{L}^{(\pm)}(E)&=A_{L}^{(\pm)}\sin\left(\frac{2\pi LE}{\hbar\omega(1\mp\mathring{\kappa})}-\frac{\pi}{2}L\sigma_{\pm}\right)\\ &\quad+B_{L}^{(\pm)}\cos\left(\frac{2\pi LE}{\hbar\omega(1\mp\mathring{\kappa})}-\frac{\pi}{2}L\sigma_{\pm}\right),\end{split}$
$\delta g_{N}^{(0)}(E)=B_{N}^{(0)}\cos\left(\frac{2\pi NE}{\hbar\omega}-\frac{\pi}{2}N\sigma_{0}\right).$
Thus, the terms $\delta g_{L}^{(\pm)}$ and $\delta g_{N}$ can be interpreted as the contributions of POs with the primitive periods
$T_{\pm}=\dfrac{2\pi}{\omega(1\mp\mathring{\kappa})}=\frac{2\pi}{\omega\mp\kappa s}\quad\mbox{and}\quad T_{0}=\dfrac{2\pi}{\omega},$
respectively. In the coupled-channel WKB prescription, two adiabatic Hamiltonians $H_{\pm}=H_{0}\mp\kappa|{\bm{r}}\times{\bm{p}}|s$ are obtained, and circular PO families with periods $T=T_{+}$ and $T_{-}$ are found in the Hamiltonians $H_{+}$ and $H_{-}$, respectively. There is also a diametric PO family with $T=T_{0}$, but it has the orbital angular momentum $\bm{l}={\bm{r}}\times{\bm{p}}=0$ and lies on the mode conversion point $(H_{+}=H_{-})$. Hence, its contribution cannot be evaluated in this prescription. Another problem is the inconsistency in the semiclassical order of the orbit contributions. In general, the amplitude factor of $D$ parameter family is proportional to $\hbar^{-D/2}$. Since both the circular and diametric orbit families have $D=2$ due to the spherical symmetry, the amplitude is expected to be proportional to $\hbar^{-1}$, however, the leading-order term of the amplitudes $A_{L}^{(\pm)}$ in the Berry-Tabor formula (6) is proportional to 2superscriptPlanck-constant-over-$\hbar^{-2}$, which suggests a contribution of 4-parameter family. Then, it is an interesting question if the classical POs in which spin dynamics is explicitly considered in the EOM could help us overcome the above difficulties and obtain a more appropriate quantum-classical correspondence.
This paper is organized as follows. Sec. II recapitulates the basic concepts of the classical counterpart of the spin and its dynamics derived by the spin coherent-state path integral representation. In Sec. III, special types of classical orbits with frozen spin and their properties are discussed. In Sec. IV, generic solutions to the classical EOM with spin precession are derived. Using this solution, periodic orbits are investigated and their bifurcation scenarios with varying energies are clarified. Sec. V is devoted to a summary and perspective.
II Spin canonical variables and the classical equations of motion
The spin is a quantum mechanical quantity analogous to the angular momentum. In quantum mechanics, the spin operator $\hat{s}_{i}$ satisfies the following angular momentum algebra,
$[\hat{s}_{i},\hat{s}_{j}]=i\hbar\varepsilon_{ijk}\hat{s_{k}},$ (7)
where $\varepsilon_{ijk}$ stands for a completely antisymmetric unit tensor (Levi-Civita symbol), and the sum is taken over identical subscripts in each term following the usual Einstein convention. Spin raising and lowering operators $\hat{s}_{\pm}\equiv\hat{s}_{x}\pm i\hat{s}_{y}$ satisfy
$[\hat{s}_{+},\hat{s}_{-}]=2\hbar\hat{s}_{z},\quad[\hat{s}_{z},\hat{s}_{\pm}]=\pm\hbar\hat{s}_{\pm}.$ (8)
The classical counterpart of the spin is introduced via the spin coherent state[42, 37] defined by
$|\zeta\rangle=\frac{1}{(1+|\zeta|^{2})^{s}}e^{\zeta\hat{s}_{+}/\hbar}|\mathring{s},-\mathring{s}\rangle,$ (9)
where $\mathring{s}$ represents the spin quantum number and $|\mathring{s},m_{s}\rangle$ represents the spin eigenstate satisfying
$\hat{s}^{2}|\mathring{s},m_{s}\rangle=\hbar^{2}\mathring{s}(\mathring{s}+1)|\mathring{s},m_{s}\rangle$
$\hat{s}_{z}|\mathring{s},m_{s}\rangle=\hbar m_{s}|\mathring{s},m_{s}\rangle,\quad m_{s}=-\mathring{s},-\mathring{s}+1,\cdots,\mathring{s}.$ (10)
$\zeta$ is an arbitrary complex number, and $|\zeta\rangle$ forms a non-orthogonal, overcomplete set of spin state vectors. The classical counterparts of the spin variables are defined through the expectation values of the spin operators as
$\begin{split}s_{x}&=\langle\zeta|\hat{s}_{x}|\zeta\rangle=s\sin\vartheta\cos\varphi,\\ s_{y}&=\langle\zeta|\hat{s}_{y}|\zeta\rangle=s\sin\vartheta\sin\varphi,\\ s_{z}&=\langle\zeta|\hat{s}_{z}|\zeta\rangle=s\cos\vartheta,\end{split}$ (11)
where the angles $\vartheta,\varphi$ are related to the complex parameter $\zeta$ by
$\frac{1}{\zeta}=\tan\frac{\vartheta}{2}e^{i\varphi}.$ (12)
The semiclassical approximation to the coherent state path integral representation[37] derives the classical EOM as
$\dot{\zeta}=\frac{(1+|\zeta|^{2})^{2}}{2is}\frac{\partial{\mathscr{H}}}{\partial\zeta^{*}},\quad\dot{\zeta}^{*}=-\frac{(1+|\zeta|^{2})^{2}}{2is}\frac{\partial{\mathscr{H}}}{\partial\zeta},$ (13)
with the Hamiltonian ${\mathscr{H}}=\langle\zeta|\hat{H}|\zeta\rangle$. The EOM are translated to those for the spin variables as
$\dot{s}_{z}=-\frac{\partial{\mathscr{H}}}{\partial\varphi},\quad\dot{\varphi}=\frac{\partial{\mathscr{H}}}{\partial s_{z}},$ (14)
showing that the classical spin motion can be described by a set of canonical variables $(q_{s},p_{s})=(\varphi,s_{z})$. Then, the Poisson brackets between spin components satisfy
$\{s_{i},s_{j}\}=\frac{\partial s_{i}}{\partial q_{s}}\frac{\partial s_{j}}{\partial p_{s}}-\frac{\partial s_{i}}{\partial p_{s}}\frac{\partial s_{j}}{\partial q_{s}}=\varepsilon_{ijk}s_{k}$ (15)
which is in exact correspondence with the quantum spin commutation relation (7).
For the Hamiltonian (2), classical EOM are expressed as
$\dot{{\bm{r}}}=\frac{\partial H}{\partial{\bm{p}}}=\frac{{\bm{p}}}{M}-\kappa{\bm{s}}\times{\bm{r}},$ (16a)
$\dot{{\bm{p}}}=-\frac{\partial H}{\partial{\bm{r}}}=-M\omega^{2}{\bm{r}}+\kappa{\bm{p}}\times{\bm{s}},$ (16b)
$\dot{{\bm{s}}}=\{{\bm{s}},H\}=-\kappa({\bm{r}}\times{\bm{p}})\times{\bm{s}}.$ (16c)
Time evolution of the orbital angular momentum ${\bm{l}}={\bm{r}}\times{\bm{p}}$ becomes
$\dot{{\bm{l}}}=\kappa({\bm{r}}\times{\bm{p}})\times{\bm{s}}$ (17)
and the total angular momentum ${\bm{j}}$ is shown to be conserved:
$\dot{{\bm{l}}}+\dot{{\bm{s}}}=0,\quad{\bm{l}}+{\bm{s}}={\bm{j}}=\mbox{const.}$. (18)
Eqs. (16c) and (17) can be rewritten as
$\dot{{\bm{s}}}=-\kappa{\bm{j}}\times{\bm{s}},\quad\dot{{\bm{l}}}=-\kappa{\bm{s}}\times{\bm{l}}=-\kappa{\bm{j}}\times{\bm{l}}$ (19)
and one sees that both ${\bm{s}}$ and ${\bm{l}}$ precess around the conserved vector ${\bm{j}}$ with angular velocity $\bm{\omega}_{p}=-\kappa{\bm{j}}$ (see Fig. 1).
Figure 1: Precession of angular momentum vectors.
III Orbits with frozen spin
By putting ${\bm{j}}=(0,0,j)$, the spin components satisfying Eq. (19) are explicitly given by the functions of time $t$ as
$s_{x}=s_{\perp}\sin\omega_{p}t,\quad s_{y}=s_{\perp}\cos\omega_{p}t,$ (20)
$(s_{\perp}=\sqrt{s^{2}-s_{z}^{2}},\quad\omega_{p}=\kappa j)$
with $s_{z}$ being constant. First, let us consider the orbit with $s_{\perp}=0$, for which spin vector ${\bm{s}}=(0,0,s_{z})$ is conserved. In this case, the EOM become
$\dot{x}=\frac{p_{x}}{M}+\kappa s_{z}y,\quad\dot{p}_{x}=-M\omega^{2}x+\kappa s_{z}p_{y},$
$\dot{y}=\frac{p_{y}}{M}-\kappa s_{z}x,\quad\dot{p}_{y}=-M\omega^{2}y-\kappa s_{z}p_{x},$
$\dot{z}=\frac{p_{z}}{M},\quad\dot{p}_{z}=-M\omega^{2}z$ (21)
with $l_{x}=l_{y}=0$. The motion in $z$ direction becomes a harmonic vibration with angular frequency $\omega$. Taking $x=y=0$, one obtains the diametric PO
$z(t)=z_{0}\sin\omega t,\quad z_{0}=\sqrt{\frac{2E}{M\omega^{2}}}$ (22)
with period $T=2\pi/\omega$. Since the orientation of the spin (the choice of $z$ axis) is arbitrary, the orbit forms a 2-parameter family under given energy $E$.
Figure 2: Circular families of frozen-spin orbits.
On the other hand, one has a planar orbit in $xy$ plane by putting $z=0$. To solve the EOM (21) for $x$ and $y$ parts, it is convenient to define the complex variables
$\xi=x-\frac{i}{M\omega}p_{x},\quad\eta=y-\frac{i}{M\omega}p_{y}.$ (23)
The EOM (21) are then transformed into
$\begin{pmatrix}\dot{\xi}\\ \dot{\eta}\end{pmatrix}=i\begin{pmatrix}\omega&-i\kappa s_{z}\\ i\kappa s_{z}&\omega\end{pmatrix}\begin{pmatrix}\xi\\ \eta\end{pmatrix}$ (24)
and one obtains two eigenmodes with angular frequencies $\omega\pm\kappa s_{z}$. For the frequency $\omega_{c}=\omega-\kappa s_{z}$, one has
$\xi=i\eta=r_{c}e^{i\omega_{c}t},$
$x=\mathop{\rm Re}\xi=r_{c}\cos\omega_{c}t,\quad y=\mathop{\rm Re}\eta=r_{c}\sin\omega_{c}t,$ (25)
which gives circular POs with angular frequencies $\omega_{c}=\omega\mp\kappa s\equiv\omega_{\pm}$ corresponding to $s_{z}=\pm s$. From the EOM (21), momentum components become
$p_{x}=M(\dot{x}-\kappa s_{z}y)=-Mr_{c}\omega\sin\omega_{c}t,$
$p_{y}=M(\dot{y}+\kappa s_{z}x)=Mr_{c}\omega\cos\omega_{c}t.$
The orbital angular momentum $l_{z}$ and energy $E$ are then expressed as
$l_{z}=xp_{y}-yp_{x}=Mr_{c}^{2}\omega,$
$E=\frac{(Mr_{c}\omega)^{2}}{2M}+\frac{M\omega^{2}r_{c}^{2}}{2}-\kappa l_{z}s_{z}=Mr_{c}^{2}\omega(\omega-\kappa s_{z})=l_{z}\omega_{c}.$
and the radii of the circular orbits are given by
$r_{c}=\sqrt{\frac{l_{z}}{M\omega}}=\sqrt{\frac{E}{M\omega\omega_{\pm}}}\equiv r_{\pm}.$ (26)
The angular frequencies $\omega_{+}$ and $\omega_{-}$ are of the POs whose orbital angular momenta are parallel and anti-parallel to the spin, respectively (see Fig. 2). Each of them forms a 2-parameter family, as well as the diametric family.
Figure 3: Examples of 3D frozen-spin orbits for (a) $\kappa s=\omega/6$ and (b) $\kappa s=\omega/4$.
In special cases where $\omega_{+}$ and $\omega_{-}$ are commensurable with each other, all planar orbits become periodic. For a spin-orbit parameter satisfying
$\kappa s=\frac{m}{n}\omega,$ (27)
with integers $n$ and $m$ being relatively prime ($n>|m|\geq 1$), period of the orbit becomes
$T=\frac{2\pi(n-m)}{\omega_{+}}=\frac{2\pi(n+m)}{\omega_{-}}=\frac{2\pi n}{\omega}.$ (28)
In this case, $\omega_{\pm}$ is also commensurable with $\omega$, and three-dimensional (3D) POs can be realized by combining motion in the $z$ direction. Such orbits have ${\bm{l}}=0$, for which the spin vector ${\bm{s}}={\bm{j}}$ is conserved, and the position vector ${\bm{r}}$ always pass through the origin (see Fig. 3).
IV Orbits with spin precession
IV.1 Bifurcations of circular orbits
If the spin direction in each circular PO is slightly tilted, the spin vector begins to precess around the total angular momentum following the EOM (19). Using Eq. (24) and the EOM (21), the orbital angular momentum of the circular PO are obtained as
$l_{z}=Mr_{\pm}^{2}\omega=\frac{E}{\omega_{\pm}}$ (29)
and the angular frequency of the spin precession is given by
$\omega_{p}=\kappa j=\kappa\left(\frac{E}{\omega_{\pm}}\pm s\right).$ (30)
If this frequency becomes commensurable with that of the orbital motion, $\omega_{\pm}$, a new PO will emerge in which orbital motion and spin precession are combined. The condition for such bifurcation is given by
$\omega_{p}=\frac{m}{k}\omega_{\pm},$ (31)
where $m$ and $k$ denote mutually prime integers. These conditions are successively satisfied by each circular PO with varying energy $E$. From Eqs. (30) and (31), the energies of these bifurcation points are given by
$E^{\pm}_{k,m}=\frac{\omega_{\pm}(\frac{m}{k}\omega_{\pm}-\kappa s_{z})}{\kappa},\quad s_{z}=\pm s$ (32)
IV.2 General solution to the equations of motion
For a given total angular momentum ${\bm{j}}=(0,0,j)$, EOM for the spin (19) can be solved, and the spin components are explicitly given as functions of time by Eqs. (20). Using this solution, the EOM for the orbital degrees of freedom are expressed as
$\begin{array}[]{rl}\dot{x}&=\dfrac{p_{x}}{M}-\kappa(zs_{y}-ys_{z}),\\[6.0pt] \dot{y}&=\dfrac{p_{y}}{M}+\kappa(zs_{x}-xs_{z}),\\[6.0pt] \dot{z}&=\dfrac{p_{z}}{M}+\kappa(xs_{y}-ys_{x}),\\[6.0pt] \dot{p}_{x}&=-M\omega^{2}x-\kappa(p_{z}s_{y}-p_{y}s_{z})\\ \dot{p}_{y}&=-M\omega^{2}y+\kappa(p_{z}s_{x}-p_{x}s_{z}),\\ \dot{p}_{z}&=-M\omega^{2}z+\kappa(p_{x}s_{y}-p_{y}s_{x})\end{array}$ (39)
The task is to find a solution to the above coupled linear equations for ${\bm{r}}$ and ${\bm{p}}$ with time-dependent coefficients $s_{x}(t)$ and $s_{y}(t)$ given by Eq. (20). By introducing complex variables
$\xi=x-\frac{ip_{x}}{M\omega},\quad\eta=y-\frac{ip_{y}}{M\omega},\quad\zeta=z-\frac{ip_{z}}{M\omega},$ (40)
EOM (39) are transformed into a compact form as
$\begin{array}[]{rl}\dot{\xi}&=i\omega\xi-\kappa(\zeta s_{y}-\eta s_{z}),\\ \dot{\eta}&=i\omega\eta+\kappa(\zeta s_{x}-\xi s_{z}),\\ \dot{\zeta}&=i\omega\zeta+\kappa(\xi s_{y}-\eta s_{x}).\end{array}$ (44)
The energy $E$ and component $l_{z}$ of the orbital angular momentum are expressed as
$E=\frac{M\omega^{2}}{2}(|\xi|^{2}+|\eta|^{2}+|\zeta|^{2})-\kappa(l_{z}s_{z}-s_{\perp}^{2}),$ (45)
$l_{z}=xp_{y}-yp_{x}=\frac{iM\omega}{2}(\xi^{*}\eta-\xi\eta^{*}).$ (46)
Because of the relation ${\bm{j}}={\bm{l}}+{\bm{s}}=(0,0,j)$, $x$ and $y$ components of the orbital angular momentum satisfy $l_{x}=-s_{x}$ and $l_{y}=-s_{y}$. By a further transformation $(u,v)=(\xi+i\eta,\xi-i\eta)$ which replaces the trigonometric functions in the coefficients of EOM with exponential functions, one obtains the relations
$E=\frac{M\omega^{2}}{2}\left(\frac{|u|^{2}+|v|^{2}}{2}+|\zeta|^{2}\right)-\kappa(l_{z}s_{z}-s_{\perp}^{2})$ (47)
$l_{z}=\frac{M\omega}{4}(|u|^{2}-|v|^{2})$ (48)
and the EOM (44) appear as
$\begin{array}[]{rl}\dot{u}&=i(\omega-\kappa s_{z})u-\kappa s_{\perp}\zeta e^{-i\omega_{p}t}\\ \dot{v}&=i(\omega+\kappa s_{z})v-\kappa s_{\perp}\zeta e^{i\omega_{p}t}\\ \dot{\zeta}&=i\omega\zeta+\tfrac{1}{2}\kappa s_{\perp}(ue^{i\omega_{p}t}+ve^{-i\omega_{p}t})\end{array}$ (52)
Putting $ue^{i\omega_{p}t}=U$ and $ve^{-i\omega_{p}t}=V$, one finally obtains coupled EOM with constant coefficients as
$\begin{array}[]{rl}\dot{\zeta}&=i\omega\zeta+\tfrac{1}{2}\kappa s_{\perp}(U+V)\\ \dot{U}-i\omega_{p}U&=i(\omega-\kappa s_{z})U-\kappa s_{\perp}\zeta\\ \dot{V}+i\omega_{p}V&=i(\omega+\kappa s_{z})V-\kappa s_{\perp}\zeta\end{array}$ (56)
that are combined into an eigenmode equation
$\frac{d}{dt}\begin{pmatrix}\zeta\\ U\\ V\end{pmatrix}=i\mathsf{M}\begin{pmatrix}\zeta\\ U\\ V\end{pmatrix},$
$\mathsf{M}=\begin{pmatrix}\omega&-i\kappa s_{\perp}/2&-i\kappa s_{\perp}/2\\ i\kappa s_{\perp}&\omega+\omega_{p}-\kappa s_{z}&0\\ i\kappa s_{\perp}&0&\omega-\omega_{p}+\kappa s_{z}\end{pmatrix}.$ (57)
By solving the secular equation $\det(\mu-\mathsf{M})=0$, the three eigenvalues $\mu$ of the matrix $\mathsf{M}$ are obtained as
$\mu=\omega,\quad\omega\pm\kappa l,$ (58)
where I have used the relation
$(\omega_{p}-\kappa s_{z})^{2}+(\kappa s_{\perp})^{2}=(\kappa l_{z})^{2}+(-\kappa l_{\perp})^{2}=(\kappa l)^{2}.$
The generic orbit can then be expressed by a linear combination of these three modes:
$\begin{pmatrix}\zeta\\ U\\ V\end{pmatrix}=\left\{\begin{pmatrix}Z_{0}\\ U_{0}\\ V_{0}\end{pmatrix}+\begin{pmatrix}Z_{+}\\ U_{+}\\ V_{+}\end{pmatrix}e^{i\kappa lt}+\begin{pmatrix}Z_{-}\\ U_{-}\\ V_{-}\end{pmatrix}e^{-i\kappa lt}\right\}e^{i\omega t},$ (59)
with the coefficients satisfying the following relation:
$U_{0}=-V_{0}=\frac{is_{\perp}}{l_{z}}Z_{0},$
$U_{\pm}=\frac{is_{\perp}}{\pm l-l_{z}}Z_{\pm}=-\frac{\pm l+l_{z}}{is_{\perp}}Z_{\pm},$
$V_{\pm}=\frac{is_{\perp}}{\pm l+l_{z}}Z_{\pm}=-\frac{\pm l-l_{z}}{is_{\perp}}Z_{\pm}.$ (60)
In the above, $l_{z}\neq 0$ is assumed, which is satisfied by most of the orbits under consideration. (Possibile POs with $l_{z}=0$ are described in Appendix B.) By applying Eqs. (59) and (60) to Eq. (48), one obtains the relation
$l_{z}$ $=\frac{1}{4}M\omega\left(|U|^{2}-|V|^{2}\right)$
$=M\omega\left[\frac{l\,l_{z}}{s_{\perp}^{2}}(|Z_{+}|^{2}-|Z_{-}|^{2})\right.$
$\hskip 34.44434pt\left.+\frac{l}{l_{z}}\mathop{\rm Re}\Bigl{\{}(Z_{0}^{*}Z_{+}-Z_{0}Z_{-}^{*})e^{i\kappa lt}\Bigr{\}}\right].$ (61)
In order that $l_{z}$ be independent of time, one must have
$Z_{0}^{*}Z_{+}-Z_{0}Z_{-}^{*}=0.$
Nonvanishing $Z_{0}$ leads to $|Z_{+}|=|Z_{-}|$ and thus to improper result $l_{z}=0$. Therefore, $Z_{0}=0$ and one obtains
$|Z_{+}|^{2}-|Z_{-}|^{2}=\frac{s_{\perp}^{2}}{M\omega l}.$ (62)
IV.3 Bridge between circular orbits
The periodicity condition is now considered. There are three independent angular frequencies; $\omega_{p}$ for the spin precession, and $\omega\pm\kappa l$ for the orbital motion in $z$ direction. The frequencies for the motion in $x$ and $y$ directions ($\omega\pm\kappa l\pm\omega_{p}$) are given by the combinations of the above three. Let us first consider the case where one of the coefficients $Z_{\pm}$ vanishes. In such case, only one of the two frequencies $\omega\pm\kappa l$ should be considered in the periodicity condition. Taking into account the relation (62), let us put
$Z_{-}=0,\quad Z_{+}=\frac{s_{\perp}}{\sqrt{M\omega l}}e^{i\gamma}$ (63)
where $\gamma$ is an arbitrary phase parameter. The periodicity condition then requires
$\omega+\kappa l=\frac{n}{m}\omega_{p}$ (64)
where $n$ and $m$ are relatively prime integers satisfying $n>m$, and the period of the PO is given by
$T=\frac{2\pi m}{\omega_{p}}=\frac{2\pi n}{\omega+\kappa l}=\frac{2\pi(n\pm m)}{\omega+\kappa l\pm\omega_{p}}.$ (65)
Rewriting Eq. (64) in terms of $j$ and $s_{z}$ yields
$\omega+\kappa\sqrt{j^{2}-2js_{z}+s^{2}}=\frac{n}{m}\kappa j,$ (66)
with which $j$ can be expressed as a function of $s_{z}$ for each set of $(n,m)$,
$\kappa j=$ $\frac{m}{n^{2}-m^{2}}\Bigl{(}n\omega-m\kappa s_{z}$
$+\sqrt{(n\omega-m\kappa s_{z})^{2}-(n^{2}-m^{2})(\omega^{2}-\kappa^{2}s^{2})}\Bigr{)}.$ (67)
Then, one can determine $\omega_{p}=\kappa j$, $l_{z}=j-s_{z}$, $l=\sqrt{l_{z}^{2}+s_{\perp}^{2}}$, and the coefficients $U_{+}$, $V_{+}$ via Eq. (63) that are necessary to calculate the PO for each value of $s_{z}$. Using these quantities, the energy (47) of the PO is obtained as
$E=\omega\sqrt{j^{2}-2js_{z}+s^{2}}-\kappa(js_{z}-s^{2}).$ (68)
Figure 4: Plot of the energy $E$ (in units of $\hbar\omega$) vs. period $T$ (in units of $2\pi/\omega$) of the periodic orbits. The vertical lines represent the diametric orbits (dotted line) and two types of circular orbits $C_{\pm}$ (dashed lines). Open circles indicate the bifurcation points (32), and the curves connected them represent the bridge orbits $(n,m)$.
Figure 4 shows the relation between energy $E$ and period $T$ of the above PO for various sets of $(n,m)$ with varying $s_{z}$ from $s$ to $-s$. Each such PO forms a bridge between the two circular orbits, as it emerge from the circular orbit C+, increasing spin obliquity with energy and finally submerge into another circular orbit C-. It can be easily checked that the above orbit coincides the circular orbits (25) in the limits $s_{z}\to\pm s$. The bifurcation scenario of the bridge orbit is illustrated in Fig. 5.
Figure 5: Illustration of the bifurcation scenario of the bridge orbit. Lower to upper panels in energy order.
Each bridge orbit forms a 3-parameter family since the orientation of the total angular momentum ${\bm{j}}$ and the phase $\gamma$ in Eq. (63) are arbitrary. Figure 6 displays some examples of bridge orbits $(n,m)$ for several values of $s_{z}$. As the spin vector tilts, the orbital ripples develop and eventually subside toward spin flip.
Figure 6: Some bridge orbits $(n,m)$ with $n-m=1$, calculated for several values of $s_{z}$ (in units of $\hbar$) between $s$ and $-s$. The parameters $\kappa=0.1\,\omega/\hbar$ and $s=\hbar/2$ are used. Note that the vertical magnification is set approximately 10 times the horizontal one.
IV.4 Secondary bifurcation
Each bridge orbit $(n,m)$ on the way from $C_{+}$ to $C_{-}$, the period $2n\pi/(\omega+\kappa l)$ never misses matching the middle of those for $C_{\pm}$, namely, a multiple of $2\pi/\omega$. Since $\omega+\kappa l$ becomes commensurable with $\omega$ at that point, it is also commensurable with $\omega-\kappa l$ and the combination of the modes $\mu=\omega\pm\kappa l$
$\begin{pmatrix}\zeta\\ ue^{i\kappa jt}\\ ve^{-i\kappa jt}\end{pmatrix}=\left\{\begin{pmatrix}Z_{+}\\ U_{+}\\ V_{+}\end{pmatrix}e^{i\kappa lt}+\begin{pmatrix}Z_{-}\\ U_{-}\\ V_{-}\end{pmatrix}e^{-i\kappa lt}\right\}e^{i\omega t}$ (69)
provides the periodic solution. The commensurability condition
$\frac{\omega+\kappa l}{n}=\frac{\kappa j}{m}=\frac{\omega}{k}$ (70)
and trigonometric inequality $|j-l|\leq s$ lead to the following relation:
$|j-l|=\left|j-\frac{n-k}{m}j\right|=\frac{|m-(n-k)|}{m}j\leq s.$ (71)
This condition is satisfied at least once for all $(n,m)$ at $k=n-m$. For sufficiently small $j$ and large $m$ satisfying $ms/j>1$, one finds several integers $k$ which satisfy (71), but solutions other than $k=n-m$ correspond to rather long orbits with period $T>2\pi/\kappa s$ and are irrelevant to the gross shell structure.
The PO is given by Eq. (69) with coefficients satisfying Eqs. (60) and (62). By inserting these relations into Eq. (47), the energy is expressed as
$E=\frac{M\omega^{2}l^{2}}{s_{\perp}^{2}}(|Z_{+}|^{2}+|Z_{-}|^{2})+\frac{1}{2}\kappa s^{2}.$ (72)
The orbit bifurcated from the bridge $(2,1)$ are displayed in Fig. 7 for several values of energy.
Figure 7: PO bifurcated from the bridge $(n,m)=(2,1)$ at several values of energy $E$ (in units of $\hbar\omega$). $\kappa=0.1\,\omega/\hbar$, $s=\hbar/2$ are taken, and $E_{\rm bif}=10.025\,\hbar\omega$.
Using Eqs. (62) and (72), one obtains
$|Z_{\pm}|^{2}=\frac{s_{\perp}^{2}}{2M\omega^{2}l^{2}}\left\{E-\tfrac{1}{2}\kappa s^{2}\pm\omega l\right\}.$ (73)
The bifurcation energy corresponding to $|Z_{-}|=0$ is given by
$E_{\rm bif}^{(n,m,k)}=\frac{n-k}{k}\frac{\omega^{2}}{\kappa}+\frac{1}{2}\kappa s^{2}.$ (74)
and the orbit exists for $E\geq E_{\rm bif}$. Each orbit $(n,m,k)$ forms a 4-parameter family since the orientation of ${\bm{j}}$ and the arguments of the complex coefficients $Z_{\pm}$ are arbitrary, and As displayed in the $E$-$T$ plots in Fig. 8, POs with the same period $T=2\pi k/\omega$ but having different geometries $(n,m,k)$ successively emerge with increasing energy.
Figure 8: $E$-$T$ plot for the orbits bifurcated from the middle of the bridge orbits. The solid and broken lines represent the circular orbits and bridge between them, respectively. The open triangles indicate the bifurcation points of the bridges, and the straight lines on the top of them represent the new orbits. The units for $T$ and $E$ are same as in Fig. 4.
V Summary
Classical motion was investigated in the extended phase space for the system of a spherical harmonic oscillator with simple spin-orbit coupling. General periodic solutions to the equations of motion were obtained and the bifurcation scenarios were fully clarified in both generic and specific values of the spin-orbit parameter. For arbitrary value of the spin-orbit parameter, there are two-parameter families of diametric and circular POs with frozen spin. There are two types of circular families C± in which spin is parallel or antiparallel to the orbital angular momentum. These circular orbits encounter successive bifurcations with increasing energy and generate three-parameter bridge PO families. On its way from one circular PO to another, each bridge encounters another bifurcation and the four-parameter PO family emerges, which has a constant period and survives for any higher energy. For specific values of spin-orbit parameter, different types of periodic solutions are also possible, for both frozen and moving spins.
The Berry-Tabor formula for the system under consideration suggests 4-parameter family of POs with period $T_{\pm}$, but the 4-parameter families found in this work has period $T_{0}$. The enhanced contribution of $T_{\pm}$ orbits might be explained as the effect of successive bifurcations of the circular POs and may include the contribution of bridge orbits whose periods are between $T_{+}$ and $T_{-}$. Concerning to the reduced contribution of the 4-parameter family with period $T_{0}$, cancellation may occur among the orbits proliferating with increasing energy and having different geometries. The trace formula in extended phase space [38] combined with the uniform approximations may solve these problems, that is left for a future subject.
Appendix A Berry-Tabor Formula
From the eigenvalue spectrum (5), densities of states $g^{(\pm)}$ for $\mathring{\jmath}=\mathring{l}\pm 1/2$ levels are given by
$g^{(+)}(E)$ $=\sum_{n=0}^{\infty}\sum_{\mathring{l}=0}^{\infty}(2\mathring{l}+2)\delta\Bigl{(}E-\hbar\omega(2n+\mathring{l}+\tfrac{3}{2}-\mathring{\kappa}\mathring{l})\Bigr{)}=\sum_{n=0}^{\infty}\sum_{\mathring{l}^{\prime}=1}^{\infty}2\mathring{l}^{\prime}\>\delta\Bigl{(}E-\hbar\omega(2n+\mathring{l}^{\prime}+\tfrac{1}{2}-\mathring{\kappa}(\mathring{l}^{\prime}-1)\Bigr{)},$
$g^{(-)}(E)$ $=\sum_{n=0}^{\infty}\sum_{\mathring{l}=1}^{\infty}2\mathring{l}\>\delta\Bigl{(}E-\hbar\omega(2n+\mathring{l}+\tfrac{3}{2}+\mathring{\kappa}(\mathring{l}+1))\Bigr{)}$ (75)
By applying the Poisson’s sum formula of the form
[…]
to the both of the double sum in Eq. (A1) and taking one of the new sums over integers using the expansion formula of trigonometric function into fractions
$\cot z=\sum_{n=-\infty}^{\infty}\frac{1}{z-n\pi}\quad\Biggl{(}\equiv\lim_{N\to\infty}\sum_{n=-N}^{N}\frac{1}{z-n\pi}\Biggr{)},\quad\mathop{\rm cosec}\nolimits^{2}z=\sum_{n=-\infty}^{\infty}\frac{1}{(z-n\pi)^{2}},$ (77)
the Berry-Tabor formula is obtaines as
$g(E)=g^{(+)}(E)+g^{(-)}(E)=\sum_{\sigma=\pm}\bar{g}^{(\sigma)}(E)+\sum_{\sigma=\pm}\sum_{L=1}^{\infty}\delta g_{L}^{(\sigma)}(E)+\sum_{N=1}^{\infty}\delta g_{N}^{(0)}(E),$ (78)
$\bar{g}^{(\pm)}(E)=\frac{\epsilon_{\pm}^{2}}{2\hbar\omega(1\mp\mathring{\kappa})^{2}},\quad\epsilon_{\pm}=\frac{E}{\hbar\omega}-1\pm\frac{1}{2}-\mathring{\kappa}$ (79)
$\delta g_{L}^{(\pm)}(E)=\frac{2}{\hbar\omega(1\mp\mathring{\kappa})^{2}}\left(\epsilon_{\pm}\cot\frac{2\pi L}{1\mp\mathring{\kappa}}\sin\frac{2\pi L}{1\mp\mathring{\kappa}}\epsilon_{\pm}+\mathop{\rm cosec}\nolimits^{2}\frac{2\pi L}{1\mp\mathring{\kappa}}\cos\frac{2\pi L}{1\mp\mathring{\kappa}}\epsilon_{\pm}\right),$ (80)
$\delta g_{N}^{(0)}(E)=-\frac{1}{\hbar\omega}\mathop{\rm cosec}\nolimits^{2}(\mathring{\kappa}\pi N)\cos(2\pi N\epsilon_{\pm}),$ (81)
which is equivalent to that in [36]. It appears to provide the trace formula (1) with contributions of POs with primitive periods $T_{\pm}=2\pi/\omega(1\mp\mathring{\kappa})$ and $T_{0}=2\pi/\omega$, but the irrational “Maslov indices” and the amplitude factors have yet to be explained by the classical dynamics.
Appendix B Other minor orbits with spin precession
For the sake of completeness, I consider here the $l_{z}=0$ POs with spin precession, which are not considered in the main part because of their insignificant contribution to the trace formula. Note that $j=s_{z}$ and $l=s_{\perp}$, and the coefficients in Eq. (59) are related as
$Z_{0}=0,\quad U_{0}=-V_{0},\quad U_{\pm}=V_{\pm}=\pm iZ_{\pm}.$
Inserting these into the conditions
$l_{z}=\frac{M\omega}{4}(|u|^{2}-|v|^{2})=0,$
$l_{x}+il_{y}=\frac{M\omega}{2}(\zeta v^{*}-\zeta^{*}u)=-is_{\perp}e^{-i\omega_{p}t},$
$E=\frac{M\omega^{2}}{2}\left(\frac{|u|^{2}+|v|^{2}}{2}+|\zeta|^{2}\right)+\kappa s_{\perp}^{2},$
one has
$U_{0}=0,\quad|Z_{+}|^{2}-|Z_{-}|^{2}=\frac{s_{\perp}}{M\omega},\quad|Z_{+}|^{2}+|Z_{-}|^{2}=\frac{E-\kappa s_{\perp}^{2}}{M\omega^{2}}.$ (82)
For $Z_{-}=0$, the periodicity condition is
$\frac{\kappa s_{z}}{m}=\frac{\omega+\kappa s_{\perp}}{n}$
and $(s_{z},s_{\perp})$ are determined as function of $\kappa$ for given set of $(n,m)$. The period is given by
$T=\frac{2\pi n}{\omega+\kappa s_{\perp}}=\frac{2\pi m}{\kappa s_{z}},$
where $n>\omega/\kappa s$ becomes a large integer for a typical value of $\kappa$ for atomic nuclei ($\mathring{\kappa}\approx 0.06$) and corresponds to a long PO. As shown from Eq. (82), this orbit can exist only at the energy
$E=\omega s_{\perp}+\kappa s_{\perp}^{2}$
and cannot contribute to the shell structure. With varying spin-orbit parameter $\kappa$, it disappears in the limit $s_{\perp}\to 0$ ($\kappa\to m\omega/ns$).
Figure 9: $l_{z}=0$ PO with spin precession. Broken curves represent $Z_{-}~{}0$ orbit $(n,m)$ with $m=1$, and the solid lines on top of the bifurcation point marked by open circles represent $Z_{-}\neq 0$ orbit $(n,m,k)$. Energy of PO (in units of $\hbar\omega$) is plotted as a function of the LS coupling strength $\kappa$ (in units of $\omega/\hbar$). Broken curves represent the PO with $Z_{-}=0$, and the open dots represent bifurcation points. At those values of spin-orbit parameter $\kappa$, the 4 parameter PO families with $Z_{-}\neq 0$ having constant periods $T=2\pi k/\omega$ ($k<n$) emerge, and they survive for higher energies.
For $Z_{-}\neq 0$, the periodicity condition becomes
$\frac{\kappa s_{z}}{m}=\frac{\omega+\kappa s_{\perp}}{n}=\frac{\omega-\kappa s_{\perp}}{n^{\prime}}=\frac{\omega}{k}\quad(n^{\prime}=2k-n),$
$\quad\frac{s_{z}}{s_{\perp}}=\frac{m}{n-k},$
and the period is given by
$T=\frac{2\pi k}{\omega}.$
This condition is satisfied only by specific values of $\kappa$ and certain spin orientations for which $\kappa s_{z}$ and $\kappa s_{\perp}$ are both commensurable with $\omega$, but the orbit can exist for continuous values of energy with coefficients
$|Z_{\pm}|^{2}=\frac{1}{2}\left(\frac{E-\kappa s_{\perp}^{2}}{M\omega^{2}}\pm\frac{s_{\perp}}{M\omega}\right),\quad E\geq E_{\rm bif}=\omega s_{\perp}+\kappa s_{\perp}^{2}.$
$k=m\omega/\kappa s_{z}>\omega/\kappa s$ is a large integer, which again gives a long PO. Figure 9 shows where the shortest $l_{z}=0$ POs are found that exist around $\kappa=0.12\,\omega/\hbar$. In the case of $(n,m,k)=(25,1,24)$, for instance, one sees
$s_{z}=s_{\perp}=\frac{s}{\sqrt{2}},\quad\kappa=\frac{\omega}{24s_{z}}\fallingdotseq 0.118\,\frac{\omega}{\hbar},\quad E_{\rm bif}\fallingdotseq 0.368\,\hbar\omega.$

### v2

@@
 A complete analysis of classical periodic orbits (POs) and their bifurcations was conducted in spherical harmonic oscillator system with spin-orbit coupling.
-The motion of the spin is explicitly considered using the spin canonical variables derived by semiclassical approximation to the spin coherent state path integral representation.
+The motion of the spin is explicitly considered using the spin canonical variables derived by semiclassical approximation to the spin coherent state path-integral representation.
 In addition to the diametric and two circular PO families with frozen spin, solutions that bridge two circular POs are found in which orbital motion is coupled to spin precession.
-In addition, each bridge encounters a secondary bifurcation on the way from one circular PO to the other and generates a new PO, that survives at higher energies while maintaining a constant period.
+In addition, each bridge encounters a secondary bifurcation on the way from one circular PO to the other and generates a new PO that survives at higher energies while maintaining a constant period.
 The generic expressions for those POs are obtained explicitly, and all the above peculiar bifurcation scenarios are described fully analytically.
 I Introduction
-Classical periodic orbits (POs) play central roles in semiclassical quantization for both integrable and nonintegral Hamiltonian systems[1, 2, 3, 4].
-Gutzwiller derived a famous trace formula that represents the quantum density of states in terms of the contributions of classical POs[5], based on the semiclassical approximation to the path integral representation using the stationary phase method.
-The formula works paricularly well for strongly chaotic systems.
-In another approach, Balian and Bloch derived the semiclassical density of states for infinite well potential systems of arbitrary boundary shapes[6].
-It is based on the multiple reflection expansion of the transition amplitude, and can be applied to both integrable and nonintegrable models.
-For integrable Hamiltonian systems whose energy spectra can be approximately obtained by the Einstein-Brillouin-Keller quantization rule (an extension of the Bohr-Sommerfeld quantization condition to multiply-periodic systems of arbitrary dimensions), Berry and Tabor showed that the oscillating part of the level density is generally represented in the form of the periodic orbit sum[7].
-The extension of the Gutzwiller formula to the case of continuous symmetries was conducted by Creagh and Littlejohn[8, 9].
-Its extension to bifurcation regions, where the original form of the Gutzwiller formula breaks down, was proposed by Sieber et al.
-using the uniform approximations[10, 11, 12, 13] and by Magner et al.
-using the improved stationary-phase method [14, 15].
-In general, semiclassical approximation of quantum mechanical level density
-$g_{\rm qm}(E)=\sum_{n}\delta(E-E_{n})$
-for a conservative Hamiltonian system is given by a trace formula expressed in the form
-$g_{\rm cl}(E)=\bar{g}(E)+\sum_{\rm PO}A_{\rm PO}\cos\left(\tfrac{1}{\hbar}S_{\rm PO}-\tfrac{\pi}{2}\mu_{\rm PO}\right).$ (1)
-[…]
-One of the most significant advantages in the trace formula is that it can explain the origin of the regular oscillations in level distribution, which is generally cannot be understood by purely quantum mechanical concepts.
-It is especially useful for analyzing gross shell structures in finite quantum systems.
-It has been applied to nuclear deformation soon after its derivation[16, 17], and then to various quantum systems such as metallic clusters[18, 19, 20] and quantum dots[21].
-Periodic-orbit theory (POT) has also been applied to the problems of quantum transport.
-The oscillation of magnetoresistance in mesocopic conductors, which is related to the electron density of states at the Fermi energy, is successfully explained by the contributing POs to the semiclassical level density[22, 23].
-The significance of PO bifurcations as the microscopic origin of the deformed shell structure has been emphasized in [20, 24, 14, 25, 26, 27, 28], and its effect on the magnetoresistance was discussed in [21].
-In addition to the orbital degrees of freedom, quantum particles have a spin degree of freedom.
-The Dirac Hamiltonian for a charged fermion in an electromagnetic field includes the spin-orbit potential, whose effect can be measured as the spin-orbit splitting of the atomic energy levels.
-In atomic nuclei, strong spin-orbit coupling is inevitable in the mean field of a nucleon to account for the nuclear “magic numbers”, numbers of constituent particles (neutrons or protons) of closed-shell configurations[29, 30, 31].
-Thus, the inclusion of spin in semiclassical POT is important for a deeper understanding of nuclear shell structures and related phenomena.
-To deal with spin-orbit coupling in semiclassical theory, two different prescriptions have been proposed.
-The first one is the coupled-channel WKB method[32, 33] in which the multicomponent wave equation is diagonalized using the Wigner-Weyl calculus and is reduced to a set of uncoupled equations corresponding to the individual channels.
-This method was applied to the problem of the spin-orbit coupling[34] by treating the spin as a channel.
-This is a kind of adiabatic approximation in which spin is regarded as a slow variable.
-The semiclassical trace formula is then constructed in terms of the POs in those adiabatic Hamiltonians that correspond to individual spin channels.
-The weakness of this prescription lies in the occurrence of the so-called mode conversion.
-Adiabatic approximation cannot be justified in the mode conversion region where two adiabatic Hamiltonians cross with each other and classical dynamics become singular.
-To deal with this problem, a use of diabatic representation has been proposed, which allows a spin flip at the mode conversion point to remove the singularity[35].
-It seems working well in some numerical applications, however, it remains ambiguous and lacks theoretical basis[36].
+Harmonic oscillator (HO) system is one of the fundamental systems both in quantum and classical dynamics.
+Three-dimensional isotropic HO system shows peculiar simplicity due to the dynamical SU(3) symmetry.
+In quantum mechanics, it has strongly degenerate equidistant energy levels, and in classical mechanics, all the trajectories are periodic.
+This system is still integrable and exactly solvable after inclusion of spin-orbit coupling given by a simple scalar product of spin and orbital angular momentum vectors.
+Such Hamiltonian is expressed as
+$H=\frac{{\bm{p}}^{2}}{2M}+\frac{M\omega^{2}{\bm{r}}^{2}}{2}-\kappa{\bm{s}}\cdot({\bm{r}}\times{\bm{p}}).$ (1)
+Here $M$ is the particle mass, $\omega$ is the oscillator frequency, and $\kappa$ is the spin-orbit coupling strength.
+This Hamiltonian appears as a simple model for the nuclear mean field[1, 2].
+This type of spin-orbit coupling appears in the nonrelativistic approximation to the Dirac equation for spin $1/2$ charged fermions under Coulomb field.
+In the nuclear mean field, strong spin-orbit coupling is known to exist for both protons and neutrons which comes from a different origin than the case of the Dirac equation.
+It plays a significant role in explaining the nuclear “magic numbers” in more realistic mean field potentials[3, 4, 2].
+In analyzing single-particle shell structures, namely, gross structures of fluctuations in the single-particle level densities, semiclassical periodic-orbit theory (POT) [5, 6, 7, 8, 9, 10] is very useful.
+It represents the fluctuating part of the level density as the contribution of classical periodic orbits (POs) in the corresponding classical Hamiltonian system.
+This POT was applied to analyze the gross shell structures in nuclear and cluster systems[11, 12, 13, 14, 15].
+If one wants to apply POT to the system with spin degree of freedom, one has to define the classical counterpart of the spin, which is purely quantum mechanical object.
+To treat spin-orbit coupling in semiclassical POT, the Wentzel-Kramers-Brillouin (WKB) method was first proposed by Littlejohn and Flynn[16] in which the multicomponent wave equation was diagonalized using the Wigner-Weyl calculus and was reduced to a set of uncoupled equations corresponding to the individual spin channels[17, 18].
+This is a kind of adiabatic approximation which treats spin as slow variable.
+However, it suffers a serious problem of mode conversions where adiabatic approximation breaks down and the classical motion becomes singular.
+A diabatic correction at the mode conversion point has been proposed[19] which seems to be appropriate in some numerical applications, however, it remains ambiguous and lacks theoretical basis[20].
 In another prescription, the spin degree of freedom is explicitly incorporated into the equations of motion (EOM).
@@
 The classical dynamics of the spin are derived by a semiclassical approximation of the spin coherent state path integral representation[21].
-Theoretically, the semiclassical approximation is justified for large spins, but in practice, it also gives reasonable results for small spins in many cases.
-Thus, we can expect it to provide a better prescription for a semiclassical treatment of the spin, free from the problems like mode conversions.
-The compensation is the increased degrees of freedom, which usually complicate classical motions.
-Pletyukov and Zeitsef derived a trace formula which explicitly consider the spin motion[38].
-They also showed that the formula works well for small spins (such as $s=\hbar/2$ for electrons and nucleons) in case Hamiltonians depend linearly on the spin.
-Thus, the proposed method can be safely applied to the spin-orbit problem in electron and nucleon systems.
-The application to the 2-dimensional Rashba model to electrons in quantum dots was performed in [36, 39].
-In this paper, I investigate the 3-dimensional isotropic harmonic oscillator model with a simple spin-orbit coupling term:
-$H=\frac{{\bm{p}}^{2}}{2M}+\frac{M\omega^{2}{\bm{r}}^{2}}{2}-\kappa({\bm{r}}\times{\bm{p}})\cdot{\bm{s}}.$ (2)
-This Hamiltonian is physically important as the mean field model for light nuclei, corresponding to the Nilsson model without $l^{2}$ term[40, 31].
-Classical dynamics in the same Hamiltonian have been studied numerically in [41].
-Quantum energy eigenvalue problem for the Hamiltonian (2) is exactly solved, and the energy eigenvalues for particles with spin $s=\hbar/2$ are expressed as
-[…]
-$(\mathring{\kappa}=\kappa s/\omega,\quad n=0,1,2,\cdots,\quad\mathring{l}=0,1,2,\cdots)$
-where $n,\mathring{l},\mathring{\jmath}$ are radial, spin, orbital angular and total angular momentum quantum numbers, respectively.
-Each level specified by $\mathring{\jmath}$ has $(2\mathring{\jmath}+1)$-fold degeneracies because the energy eigenvalues are independent on the magnetic quantum number $m_{j}$ ($-\mathring{\jmath}\leq m_{j}\leq\mathring{\jmath}$).
-For a system in which the energy eigenvalues are explicitly given in terms of the quantum numbers, Berry-Tabor formula[7] can be applied as the semiclassical density of states, and the following formula has been obtained in Ref.
-[36] (see also Appendix A)
-$g(E)=\bar{g}(E)+\sum_{\sigma=\pm}\sum_{L=1}^{\infty}\delta g_{L}^{(\sigma)}(E)+\sum_{N=1}^{\infty}\delta g_{N}^{(0)}(E),$ (6)
-$\begin{split}\delta g_{L}^{(\pm)}(E)&=A_{L}^{(\pm)}\sin\left(\frac{2\pi LE}{\hbar\omega(1\mp\mathring{\kappa})}-\frac{\pi}{2}L\sigma_{\pm}\right)\\ &\quad+B_{L}^{(\pm)}\cos\left(\frac{2\pi LE}{\hbar\omega(1\mp\mathring{\kappa})}-\frac{\pi}{2}L\sigma_{\pm}\right),\end{split}$
-$\delta g_{N}^{(0)}(E)=B_{N}^{(0)}\cos\left(\frac{2\pi NE}{\hbar\omega}-\frac{\pi}{2}N\sigma_{0}\right).$
-Thus, the terms $\delta g_{L}^{(\pm)}$ and $\delta g_{N}$ can be interpreted as the contributions of POs with the primitive periods
-$T_{\pm}=\dfrac{2\pi}{\omega(1\mp\mathring{\kappa})}=\frac{2\pi}{\omega\mp\kappa s}\quad\mbox{and}\quad T_{0}=\dfrac{2\pi}{\omega},$
-respectively.
-In the coupled-channel WKB prescription, two adiabatic Hamiltonians $H_{\pm}=H_{0}\mp\kappa|{\bm{r}}\times{\bm{p}}|s$ are obtained, and circular PO families with periods $T=T_{+}$ and $T_{-}$ are found in the Hamiltonians $H_{+}$ and $H_{-}$, respectively.
-There is also a diametric PO family with $T=T_{0}$, but it has the orbital angular momentum $\bm{l}={\bm{r}}\times{\bm{p}}=0$ and lies on the mode conversion point $(H_{+}=H_{-})$.
-Hence, its contribution cannot be evaluated in this prescription.
-Another problem is the inconsistency in the semiclassical order of the orbit contributions.
-In general, the amplitude factor of $D$ parameter family is proportional to $\hbar^{-D/2}$.
-Since both the circular and diametric orbit families have $D=2$ due to the spherical symmetry, the amplitude is expected to be proportional to $\hbar^{-1}$, however, the leading-order term of the amplitudes $A_{L}^{(\pm)}$ in the Berry-Tabor formula (6) is proportional to 2superscriptPlanck-constant-over-$\hbar^{-2}$, which suggests a contribution of 4-parameter family.
-Then, it is an interesting question if the classical POs in which spin dynamics is explicitly considered in the EOM could help us overcome the above difficulties and obtain a more appropriate quantum-classical correspondence.
-This paper is organized as follows.
-Sec.
-II recapitulates the basic concepts of the classical counterpart of the spin and its dynamics derived by the spin coherent-state path integral representation.
+Along with this prescription, the semiclassical trace formula for system with spin degree of freedom was derived[22, 23] in which spin dynamics was explicitly considered in the classical motion.
+Although the semiclassical approximation was justified only for large spins, the trace formula was found to be applicable to the system of particles with small spins, such as electrons and nucleons, at least for the Hamiltonians which depend linearly on the spin.
+The Hamiltonian (1) certainly meets this requirement.
+This finding increases the significance of investigating classical motion, especially the POs, in the extended phase space where spin motion is taken explicitly in the EOM.
+Apart from its physical importance as the nuclear mean-field model, the Hamiltonian (1) provides us an interesting and significant example of the integrable spin-orbit-coupled system.
+The classical motions in the same Hamiltonian were investigated by Rozmej and Arview[24].
+The main purpose of their work was to specify the origin of chaotic nature in the nuclear mean-field model.
+They found that the classical motion in the anisotropic HO system, which is integrable, turns gradually chaotic with increasing spin-orbit coupling strength, and concluded that the spin-orbit coupling can be regarded as one possible source of nuclear chaoticity.
+In their paper, they also studied the classical motions in the spherical potential, only numerically by the Poincaré plot for two-dimensional orbits in the angular-momentum projected space.
+The result indicates the integrability of the classical system, but the analytical studies of the orbit have not been attempted.
+Classical-quantum correspondence in the Hamiltonian similar to Eq.
+(1) was considered by Amann et al.
+in Ref.
+[20].
+They incorporated anisotropy into the HO potential and applied the WKB method.
+A good quantum-classical correspondence was found in such an anisotropic case.
+However, in the spherical case, some POs lie in the mode conversion point and the WKB method cannot be applied.
+They also found several inconsistencies between contribution of classical POs and the exact quantum spectrum.
+Then they got interest in applying the trace formula in the extended phase space they develped, but actual application has not yet been made.
+Some details on this will be discussed later in Sec.
+V.
+In this paper, I give the general analytic solution to the EOM for Hamiltonian (1) by explicitly taking account of the spin degree of freedom using the classical spin variables.
+The EOM are nonlinear in orbital and spin variables, but the spin motion can be exactly solved in the spherical model, and the spin vector is explicitly given by the function of time.
+Making use of this solution and applying a subtle time-dependent transformation of variables, the system of equations are found to arrive at time-independent homogeneous equations which can be exactly solved.
+I also construct all the possible POs.
+With varying energy, those POs exhibit peculiar interesting bifurcation scenarios, and all those bifurcations are described analytically.
+This kind of model would be not only valuable for the study of nonlinear dynamics in integrable systems but also be suggestive for the study of nonintegrable systems.
+The paper is organized as follows.
+Section II recapitulates the basic concepts of the classical counterpart of the spin and its dynamics derived by the spin coherent-state path-integral representation.
 In Sec.
@@
 IV, generic solutions to the classical EOM with spin precession are derived.
-Using this solution, periodic orbits are investigated and their bifurcation scenarios with varying energies are clarified.
-Sec.
-V is devoted to a summary and perspective.
+Using this solution, POs are investigated and their bifurcation scenarios with varying energies are clarified.
+Section V is devoted to a summary and discussion on the application of semiclassical trace formula.
 II Spin canonical variables and the classical equations of motion
-The spin is a quantum mechanical quantity analogous to the angular momentum. In quantum mechanics, the spin operator $\hat{s}_{i}$ satisfies the following angular momentum algebra,
+Spin is a quantum mechanical quantity analogous to the angular momentum. In quantum mechanics, the spin operator $\hat{s}_{i}$ satisfies the following angular momentum algebra,
 $[\hat{s}_{i},\hat{s}_{j}]=i\hbar\varepsilon_{ijk}\hat{s_{k}},$ (2)
-where $\varepsilon_{ijk}$ stands for a completely antisymmetric unit tensor (Levi-Civita symbol), and the sum is taken over identical subscripts in each term following the usual Einstein convention.
+where $\varepsilon_{ijk}$ stands for a completely antisymmetric unit tensor (the Levi-Civita symbol), and the sum is taken over identical subscripts in each term following the usual Einstein convention.
 Spin raising and lowering operators $\hat{s}_{\pm}\equiv\hat{s}_{x}\pm i\hat{s}_{y}$ satisfy
@@
 $|\zeta\rangle=\frac{1}{(1+|\zeta|^{2})^{s}}e^{\zeta\hat{s}_{+}/\hbar}|\mathring{s},-\mathring{s}\rangle,$ (4)
-where $\mathring{s}$ represents the spin quantum number and $|\mathring{s},m_{s}\rangle$ represents the spin eigenstate satisfying
-$\hat{s}^{2}|\mathring{s},m_{s}\rangle=\hbar^{2}\mathring{s}(\mathring{s}+1)|\mathring{s},m_{s}\rangle$
+where $\mathring{s}$ represents the spin quantum number [^1] In the following, I will use symbols with circle on top of them as the dimensionless version of the variables without circle.
+For example, modulus of orbital angular momentum $l=\hbar\mathring{l}$ and spin-orbit parameter $\kappa=\mathring{\kappa}\omega/s$, etc.
+in Sec.
+V.
+and $|\mathring{s},m_{s}\rangle$ represents the spin eigenstate satisfying
+$\hat{s}^{2}|\mathring{s},m_{s}\rangle=\hbar^{2}\mathring{s}(\mathring{s}+1)|\mathring{s},m_{s}\rangle,$
 $\hat{s}_{z}|\mathring{s},m_{s}\rangle=\hbar m_{s}|\mathring{s},m_{s}\rangle,\quad m_{s}=-\mathring{s},-\mathring{s}+1,\cdots,\mathring{s}.$ (5)
-$\zeta$ is an arbitrary complex number, and $|\zeta\rangle$ forms a non-orthogonal, overcomplete set of spin state vectors.
+Here $\zeta$ is an arbitrary complex number and $|\zeta\rangle$ forms a nonorthogonal, overcomplete set of spin state vectors.
 The classical counterparts of the spin variables are defined through the expectation values of the spin operators as
@@
 Time evolution of the orbital angular momentum ${\bm{l}}={\bm{r}}\times{\bm{p}}$ becomes
-$\dot{{\bm{l}}}=\kappa({\bm{r}}\times{\bm{p}})\times{\bm{s}}$ (17)
+$\dot{{\bm{l}}}=\kappa({\bm{r}}\times{\bm{p}})\times{\bm{s}},$ (12)
 and the total angular momentum ${\bm{j}}$ is shown to be conserved:
-$\dot{{\bm{l}}}+\dot{{\bm{s}}}=0,\quad{\bm{l}}+{\bm{s}}={\bm{j}}=\mbox{const.}$. (18)
-Eqs. (16c) and (17) can be rewritten as
-$\dot{{\bm{s}}}=-\kappa{\bm{j}}\times{\bm{s}},\quad\dot{{\bm{l}}}=-\kappa{\bm{s}}\times{\bm{l}}=-\kappa{\bm{j}}\times{\bm{l}}$ (19)
+$\dot{{\bm{l}}}+\dot{{\bm{s}}}=0,\quad{\bm{l}}+{\bm{s}}={\bm{j}}=\mbox{const.}.$. . (13)
+Equations (11c) and (12) can be rewritten as
+$\dot{{\bm{s}}}=-\kappa{\bm{j}}\times{\bm{s}},\quad\dot{{\bm{l}}}=-\kappa{\bm{s}}\times{\bm{l}}=-\kappa{\bm{j}}\times{\bm{l}},$ (14)
 and one sees that both ${\bm{s}}$ and ${\bm{l}}$ precess around the conserved vector ${\bm{j}}$ with angular velocity $\bm{\omega}_{p}=-\kappa{\bm{j}}$ (see Fig. 1).
@@
 $s_{x}=s_{\perp}\sin\omega_{p}t,\quad s_{y}=s_{\perp}\cos\omega_{p}t,$ (15)
-$(s_{\perp}=\sqrt{s^{2}-s_{z}^{2}},\quad\omega_{p}=\kappa j)$
+$(s_{\perp}=\sqrt{s^{2}-s_{z}^{2}},\quad\omega_{p}=\kappa j),$
 with $s_{z}$ being constant. First, let us consider the orbit with $s_{\perp}=0$, for which spin vector ${\bm{s}}=(0,0,s_{z})$ is conserved. In this case, the EOM become
@@
 $\dot{y}=\frac{p_{y}}{M}-\kappa s_{z}x,\quad\dot{p}_{y}=-M\omega^{2}y-\kappa s_{z}p_{x},$
-$\dot{z}=\frac{p_{z}}{M},\quad\dot{p}_{z}=-M\omega^{2}z$ (21)
+$\dot{z}=\frac{p_{z}}{M},\quad\dot{p}_{z}=-M\omega^{2}z,$ (16)
 with $l_{x}=l_{y}=0$. The motion in $z$ direction becomes a harmonic vibration with angular frequency $\omega$. Taking $x=y=0$, one obtains the diametric PO
-$z(t)=z_{0}\sin\omega t,\quad z_{0}=\sqrt{\frac{2E}{M\omega^{2}}}$ (22)
-with period $T=2\pi/\omega$. Since the orientation of the spin (the choice of $z$ axis) is arbitrary, the orbit forms a 2-parameter family under given energy $E$.
+$z(t)=z_{0}\sin\omega t,\quad z_{0}=\sqrt{\frac{2E}{M\omega^{2}}},$ (17)
+with period $T=2\pi/\omega$. Since the orientation of the spin (the choice of $z$ axis) is arbitrary, the orbit forms a two-parameter family under given energy $E$.
 Figure 2: Circular families of frozen-spin orbits.
-On the other hand, one has a planar orbit in $xy$ plane by putting $z=0$. To solve the EOM (21) for $x$ and $y$ parts, it is convenient to define the complex variables
+Alternatively, putting $z=0$ yields planar orbits in $xy$ plane. To solve the EOM (16) for $x$ and $y$ parts, it is convenient to define the complex variables
 $\xi=x-\frac{i}{M\omega}p_{x},\quad\eta=y-\frac{i}{M\omega}p_{y}.$ (18)
@@
 $l_{z}=xp_{y}-yp_{x}=Mr_{c}^{2}\omega,$
-$E=\frac{(Mr_{c}\omega)^{2}}{2M}+\frac{M\omega^{2}r_{c}^{2}}{2}-\kappa l_{z}s_{z}=Mr_{c}^{2}\omega(\omega-\kappa s_{z})=l_{z}\omega_{c}.$
+$E=\frac{(Mr_{c}\omega)^{2}}{2M}+\frac{M\omega^{2}r_{c}^{2}}{2}-\kappa l_{z}s_{z}=Mr_{c}^{2}\omega(\omega-\kappa s_{z})=l_{z}\omega_{c},$
 and the radii of the circular orbits are given by
 $r_{c}=\sqrt{\frac{l_{z}}{M\omega}}=\sqrt{\frac{E}{M\omega\omega_{\pm}}}\equiv r_{\pm}.$ (21)
-The angular frequencies $\omega_{+}$ and $\omega_{-}$ are of the POs whose orbital angular momenta are parallel and anti-parallel to the spin, respectively (see Fig.
+The angular frequencies $\omega_{+}$ and $\omega_{-}$ are of the POs whose orbital angular momenta are parallel and antiparallel to the spin, respectively (see Fig.
 2).
-Each of them forms a 2-parameter family, as well as the diametric family.
+Each of them forms a two-parameter family, as well as the diametric family.
 Figure 3: Examples of 3D frozen-spin orbits for (a) $\kappa s=\omega/6$ and (b) $\kappa s=\omega/4$.
-In special cases where $\omega_{+}$ and $\omega_{-}$ are commensurable with each other, all planar orbits become periodic. For a spin-orbit parameter satisfying
+In special cases where $\omega_{+}$ and $\omega_{-}$ are commensurable with each other, all planar orbits become perisodic. For a spin-orbit parameter satisfying
 $\kappa s=\frac{m}{n}\omega,$ (22)
-with integers $n$ and $m$ being relatively prime ($n>|m|\geq 1$), period of the orbit becomes
+with integers $n$ and $m$ being relatively prime ($n>|m|\geq 1$), the period of the orbit becomes
 $T=\frac{2\pi(n-m)}{\omega_{+}}=\frac{2\pi(n+m)}{\omega_{-}}=\frac{2\pi n}{\omega}.$ (23)
@@
 (25) and (26), the energies of these bifurcation points are given by
-$E^{\pm}_{k,m}=\frac{\omega_{\pm}(\frac{m}{k}\omega_{\pm}-\kappa s_{z})}{\kappa},\quad s_{z}=\pm s$ (32)
+$E^{\pm}_{k,m}=\frac{\omega_{\pm}(\frac{m}{k}\omega_{\pm}-\kappa s_{z})}{\kappa},\quad s_{z}=\pm s.$ (27)
 IV.2 General solution to the equations of motion
-For a given total angular momentum ${\bm{j}}=(0,0,j)$, EOM for the spin (19) can be solved, and the spin components are explicitly given as functions of time by Eqs.
+For a given total angular momentum ${\bm{j}}=(0,0,j)$, the EOM for the spin (14) can be solved, and the spin components are explicitly given as functions of time by Eqs.
 (15).
 Using this solution, the EOM for the orbital degrees of freedom are expressed as
-$\begin{array}[]{rl}\dot{x}&=\dfrac{p_{x}}{M}-\kappa(zs_{y}-ys_{z}),\\[6.0pt] \dot{y}&=\dfrac{p_{y}}{M}+\kappa(zs_{x}-xs_{z}),\\[6.0pt] \dot{z}&=\dfrac{p_{z}}{M}+\kappa(xs_{y}-ys_{x}),\\[6.0pt] \dot{p}_{x}&=-M\omega^{2}x-\kappa(p_{z}s_{y}-p_{y}s_{z})\\ \dot{p}_{y}&=-M\omega^{2}y+\kappa(p_{z}s_{x}-p_{x}s_{z}),\\ \dot{p}_{z}&=-M\omega^{2}z+\kappa(p_{x}s_{y}-p_{y}s_{x})\end{array}$ (39)
+$\begin{array}[]{rl}\dot{x}&=\dfrac{p_{x}}{M}-\kappa(zs_{y}-ys_{z}),\\[6.0pt] \dot{y}&=\dfrac{p_{y}}{M}+\kappa(zs_{x}-xs_{z}),\\[6.0pt] \dot{z}&=\dfrac{p_{z}}{M}+\kappa(xs_{y}-ys_{x}),\\[6.0pt] \dot{p}_{x}&=-M\omega^{2}x-\kappa(p_{z}s_{y}-p_{y}s_{z}),\\ \dot{p}_{y}&=-M\omega^{2}y+\kappa(p_{z}s_{x}-p_{x}s_{z}),\\ \dot{p}_{z}&=-M\omega^{2}z+\kappa(p_{x}s_{y}-p_{y}s_{x})\end{array}$ (34)
 The task is to find a solution to the above coupled linear equations for ${\bm{r}}$ and ${\bm{p}}$ with time-dependent coefficients $s_{x}(t)$ and $s_{y}(t)$ given by Eq.
@@
 $\xi=x-\frac{ip_{x}}{M\omega},\quad\eta=y-\frac{ip_{y}}{M\omega},\quad\zeta=z-\frac{ip_{z}}{M\omega},$ (35)
-EOM (39) are transformed into a compact form as
+the EOM (34) are transformed into a compact form as
 $\begin{array}[]{rl}\dot{\xi}&=i\omega\xi-\kappa(\zeta s_{y}-\eta s_{z}),\\ \dot{\eta}&=i\omega\eta+\kappa(\zeta s_{x}-\xi s_{z}),\\ \dot{\zeta}&=i\omega\zeta+\kappa(\xi s_{y}-\eta s_{x}).\end{array}$ (39)
@@
 $l_{z}=xp_{y}-yp_{x}=\frac{iM\omega}{2}(\xi^{*}\eta-\xi\eta^{*}).$ (41)
-Because of the relation ${\bm{j}}={\bm{l}}+{\bm{s}}=(0,0,j)$, $x$ and $y$ components of the orbital angular momentum satisfy $l_{x}=-s_{x}$ and $l_{y}=-s_{y}$.
+Because of the relation ${\bm{j}}={\bm{l}}+{\bm{s}}=(0,0,j)$, the $x$ and $y$ components of the orbital angular momentum satisfy $l_{x}=-s_{x}$ and $l_{y}=-s_{y}$.
 By a further transformation $(u,v)=(\xi+i\eta,\xi-i\eta)$ which replaces the trigonometric functions in the coefficients of EOM with exponential functions, one obtains the relations
-$E=\frac{M\omega^{2}}{2}\left(\frac{|u|^{2}+|v|^{2}}{2}+|\zeta|^{2}\right)-\kappa(l_{z}s_{z}-s_{\perp}^{2})$ (47)
-$l_{z}=\frac{M\omega}{4}(|u|^{2}-|v|^{2})$ (48)
+$E=\frac{M\omega^{2}}{2}\left(\frac{|u|^{2}+|v|^{2}}{2}+|\zeta|^{2}\right)-\kappa(l_{z}s_{z}-s_{\perp}^{2}),$ (42)
+$l_{z}=\frac{M\omega}{4}(|u|^{2}-|v|^{2}),$ (43)
 and the EOM (39) appear as
-$\begin{array}[]{rl}\dot{u}&=i(\omega-\kappa s_{z})u-\kappa s_{\perp}\zeta e^{-i\omega_{p}t}\\ \dot{v}&=i(\omega+\kappa s_{z})v-\kappa s_{\perp}\zeta e^{i\omega_{p}t}\\ \dot{\zeta}&=i\omega\zeta+\tfrac{1}{2}\kappa s_{\perp}(ue^{i\omega_{p}t}+ve^{-i\omega_{p}t})\end{array}$ (52)
+$\begin{array}[]{rl}\dot{u}&=i(\omega-\kappa s_{z})u-\kappa s_{\perp}\zeta e^{-i\omega_{p}t},\\ \dot{v}&=i(\omega+\kappa s_{z})v-\kappa s_{\perp}\zeta e^{i\omega_{p}t},\\ \dot{\zeta}&=i\omega\zeta+\tfrac{1}{2}\kappa s_{\perp}(ue^{i\omega_{p}t}+ve^{-i\omega_{p}t}).\end{array}$ (47)
 Putting $ue^{i\omega_{p}t}=U$ and $ve^{-i\omega_{p}t}=V$, one finally obtains coupled EOM with constant coefficients as
-$\begin{array}[]{rl}\dot{\zeta}&=i\omega\zeta+\tfrac{1}{2}\kappa s_{\perp}(U+V)\\ \dot{U}-i\omega_{p}U&=i(\omega-\kappa s_{z})U-\kappa s_{\perp}\zeta\\ \dot{V}+i\omega_{p}V&=i(\omega+\kappa s_{z})V-\kappa s_{\perp}\zeta\end{array}$ (56)
-that are combined into an eigenmode equation
+$\begin{array}[]{rl}\dot{\zeta}&=i\omega\zeta+\tfrac{1}{2}\kappa s_{\perp}(U+V),\\ \dot{U}-i\omega_{p}U&=i(\omega-\kappa s_{z})U-\kappa s_{\perp}\zeta,\\ \dot{V}+i\omega_{p}V&=i(\omega+\kappa s_{z})V-\kappa s_{\perp}\zeta,\end{array}$ (51)
+that are combined into a normal mode equation
 $\frac{d}{dt}\begin{pmatrix}\zeta\\ U\\ V\end{pmatrix}=i\mathsf{M}\begin{pmatrix}\zeta\\ U\\ V\end{pmatrix},$
@@
 $\mu=\omega,\quad\omega\pm\kappa l,$ (53)
-where I have used the relation
+where I use the relation
 $(\omega_{p}-\kappa s_{z})^{2}+(\kappa s_{\perp})^{2}=(\kappa l_{z})^{2}+(-\kappa l_{\perp})^{2}=(\kappa l)^{2}.$
@@
 In the above, $l_{z}\neq 0$ is assumed, which is satisfied by most of the orbits under consideration.
-(Possibile POs with $l_{z}=0$ are described in Appendix B.) By applying Eqs.
+(Possible POs with $l_{z}=0$ are described in the Appendix.) By applying Eqs.
 (54) and (55) to Eq.
@@
 $Z_{0}^{*}Z_{+}-Z_{0}Z_{-}^{*}=0.$
-Nonvanishing $Z_{0}$ leads to $|Z_{+}|=|Z_{-}|$ and thus to improper result $l_{z}=0$. Therefore, $Z_{0}=0$ and one obtains
+Nonvanishing $Z_{0}$ leads to $|Z_{+}|=|Z_{-}|$ and thus to the improper result $l_{z}=0$. Therefore, $Z_{0}=0$ and one obtains
 $|Z_{+}|^{2}-|Z_{-}|^{2}=\frac{s_{\perp}^{2}}{M\omega l}.$ (57)
@@
 The periodicity condition is now considered.
-There are three independent angular frequencies; $\omega_{p}$ for the spin precession, and $\omega\pm\kappa l$ for the orbital motion in $z$ direction.
+There are three independent angular frequencies; $\omega_{p}$ for the spin precession, and $\omega\pm\kappa l$ for the orbital motion in the $z$ direction.
 The frequencies for the motion in $x$ and $y$ directions ($\omega\pm\kappa l\pm\omega_{p}$) are given by the combinations of the above three.
 Let us first consider the case where one of the coefficients $Z_{\pm}$ vanishes.
-In such case, only one of the two frequencies $\omega\pm\kappa l$ should be considered in the periodicity condition.
+In such a case, only one of the two frequencies $\omega\pm\kappa l$ should be considered in the periodicity condition.
 Taking into account the relation (57), let us put
-$Z_{-}=0,\quad Z_{+}=\frac{s_{\perp}}{\sqrt{M\omega l}}e^{i\gamma}$ (63)
+$Z_{-}=0,\quad Z_{+}=\frac{s_{\perp}}{\sqrt{M\omega l}}e^{i\gamma},$ (58)
 where $\gamma$ is an arbitrary phase parameter. The periodicity condition then requires
-$\omega+\kappa l=\frac{n}{m}\omega_{p}$ (64)
+$\omega+\kappa l=\frac{n}{m}\omega_{p},$ (59)
 where $n$ and $m$ are relatively prime integers satisfying $n>m$, and the period of the PO is given by
@@
 Figure 4: Plot of the energy $E$ (in units of $\hbar\omega$) vs.
-period $T$ (in units of $2\pi/\omega$) of the periodic orbits.
+period $T$ (in units of $2\pi/\omega$) of the POs.
 The vertical lines represent the diametric orbits (dotted line) and two types of circular orbits $C_{\pm}$ (dashed lines).
@@
 Figure 5: Illustration of the bifurcation scenario of the bridge orbit. Lower to upper panels in energy order.
-Each bridge orbit forms a 3-parameter family since the orientation of the total angular momentum ${\bm{j}}$ and the phase $\gamma$ in Eq.
+Each bridge orbit forms a three-parameter family since the orientation of the total angular momentum ${\bm{j}}$ and the phase $\gamma$ in Eq.
 (58) are arbitrary.
 Figure 6 displays some examples of bridge orbits $(n,m)$ for several values of $s_{z}$.
-As the spin vector tilts, the orbital ripples develop and eventually subside toward spin flip.
+As the spin vector tilts, the orbital ripples develop and eventually subside toward the spin flip.
 Figure 6: Some bridge orbits $(n,m)$ with $n-m=1$, calculated for several values of $s_{z}$ (in units of $\hbar$) between $s$ and $-s$.
 The parameters $\kappa=0.1\,\omega/\hbar$ and $s=\hbar/2$ are used.
-Note that the vertical magnification is set approximately 10 times the horizontal one.
+Note that the vertical magnification is set approximately ten times the horizontal one.
 IV.4 Secondary bifurcation
-Each bridge orbit $(n,m)$ on the way from $C_{+}$ to $C_{-}$, the period $2n\pi/(\omega+\kappa l)$ never misses matching the middle of those for $C_{\pm}$, namely, a multiple of $2\pi/\omega$.
+The period of each bridge orbit $(n,m)$ always coincides with a multiple of $2\pi/\omega$ somewhere on the way from $C_{+}$ to $C_{-}$ with increasing energy.
 Since $\omega+\kappa l$ becomes commensurable with $\omega$ at that point, it is also commensurable with $\omega-\kappa l$ and the combination of the modes $\mu=\omega\pm\kappa l$
@@
 $E=\frac{M\omega^{2}l^{2}}{s_{\perp}^{2}}(|Z_{+}|^{2}+|Z_{-}|^{2})+\frac{1}{2}\kappa s^{2}.$ (67)
-The orbit bifurcated from the bridge $(2,1)$ are displayed in Fig. 7 for several values of energy.
+The orbit bifurcated from the bridge $(2,1)$ is displayed in Fig. 7 for several values of energy.
 Figure 7: PO bifurcated from the bridge $(n,m)=(2,1)$ at several values of energy $E$ (in units of $\hbar\omega$).
@@
 The bifurcation energy corresponding to $|Z_{-}|=0$ is given by
-$E_{\rm bif}^{(n,m,k)}=\frac{n-k}{k}\frac{\omega^{2}}{\kappa}+\frac{1}{2}\kappa s^{2}.$ (74)
+$E_{\rm bif}^{(n,m,k)}=\frac{n-k}{k}\frac{\omega^{2}}{\kappa}+\frac{1}{2}\kappa s^{2},$ (69)
 and the orbit exists for $E\geq E_{\rm bif}$.
-Each orbit $(n,m,k)$ forms a 4-parameter family since the orientation of ${\bm{j}}$ and the arguments of the complex coefficients $Z_{\pm}$ are arbitrary, and As displayed in the $E$-$T$ plots in Fig.
+Each orbit $(n,m,k)$ forms a four-parameter family since the orientation of ${\bm{j}}$ and the arguments $\gamma_{\pm}$ of the complex coefficients $Z_{\pm}=|Z_{\pm}|e^{i\gamma_{\pm}}$ are arbitrary.
+As displayed in the $E$-$T$ plots in Fig.
 8, POs with the same period $T=2\pi k/\omega$ but having different geometries $(n,m,k)$ successively emerge with increasing energy.
@@
 4.
-V Summary
+V Summary and discussion
 Classical motion was investigated in the extended phase space for the system of a spherical harmonic oscillator with simple spin-orbit coupling.
-General periodic solutions to the equations of motion were obtained and the bifurcation scenarios were fully clarified in both generic and specific values of the spin-orbit parameter.
-For arbitrary value of the spin-orbit parameter, there are two-parameter families of diametric and circular POs with frozen spin.
-There are two types of circular families C± in which spin is parallel or antiparallel to the orbital angular momentum.
+General solutions to the EOM and all the POs are obtained analytically, and the bifurcation scenarios are fully clarified in both generic and specific values of the spin-orbit parameter.
+For arbitrary value of the spin-orbit parameter, there are two kinds of degenerate two-parameter families of POs, diametric and circular ones, with frozen spin.
+There are two circular families C± in which spin is parallel or antiparallel to the orbital angular momentum.
 These circular orbits encounter successive bifurcations with increasing energy and generate three-parameter bridge PO families.
@@
 For specific values of spin-orbit parameter, different types of periodic solutions are also possible, for both frozen and moving spins.
-The Berry-Tabor formula for the system under consideration suggests 4-parameter family of POs with period $T_{\pm}$, but the 4-parameter families found in this work has period $T_{0}$.
-The enhanced contribution of $T_{\pm}$ orbits might be explained as the effect of successive bifurcations of the circular POs and may include the contribution of bridge orbits whose periods are between $T_{+}$ and $T_{-}$.
-Concerning to the reduced contribution of the 4-parameter family with period $T_{0}$, cancellation may occur among the orbits proliferating with increasing energy and having different geometries.
-The trace formula in extended phase space [38] combined with the uniform approximations may solve these problems, that is left for a future subject.
-Appendix A Berry-Tabor Formula
-From the eigenvalue spectrum (5), densities of states $g^{(\pm)}$ for $\mathring{\jmath}=\mathring{l}\pm 1/2$ levels are given by
-$g^{(+)}(E)$ $=\sum_{n=0}^{\infty}\sum_{\mathring{l}=0}^{\infty}(2\mathring{l}+2)\delta\Bigl{(}E-\hbar\omega(2n+\mathring{l}+\tfrac{3}{2}-\mathring{\kappa}\mathring{l})\Bigr{)}=\sum_{n=0}^{\infty}\sum_{\mathring{l}^{\prime}=1}^{\infty}2\mathring{l}^{\prime}\>\delta\Bigl{(}E-\hbar\omega(2n+\mathring{l}^{\prime}+\tfrac{1}{2}-\mathring{\kappa}(\mathring{l}^{\prime}-1)\Bigr{)},$
-$g^{(-)}(E)$ $=\sum_{n=0}^{\infty}\sum_{\mathring{l}=1}^{\infty}2\mathring{l}\>\delta\Bigl{(}E-\hbar\omega(2n+\mathring{l}+\tfrac{3}{2}+\mathring{\kappa}(\mathring{l}+1))\Bigr{)}$ (75)
-By applying the Poisson’s sum formula of the form
+Classical-quantum correspondence for this system was considered in Ref.
+[20].
+In the direction from quantum to classical, the authors constructed the exact trace formula analogous to the Berry-Tabor formula for multiply periodic systems[26].
+Quantum energy eigenvalues are given explicitly in terms of the quantum numbers as
 […]
-to the both of the double sum in Eq. (A1) and taking one of the new sums over integers using the expansion formula of trigonometric function into fractions
-$\cot z=\sum_{n=-\infty}^{\infty}\frac{1}{z-n\pi}\quad\Biggl{(}\equiv\lim_{N\to\infty}\sum_{n=-N}^{N}\frac{1}{z-n\pi}\Biggr{)},\quad\mathop{\rm cosec}\nolimits^{2}z=\sum_{n=-\infty}^{\infty}\frac{1}{(z-n\pi)^{2}},$ (77)
-the Berry-Tabor formula is obtaines as
-$g(E)=g^{(+)}(E)+g^{(-)}(E)=\sum_{\sigma=\pm}\bar{g}^{(\sigma)}(E)+\sum_{\sigma=\pm}\sum_{L=1}^{\infty}\delta g_{L}^{(\sigma)}(E)+\sum_{N=1}^{\infty}\delta g_{N}^{(0)}(E),$ (78)
-$\bar{g}^{(\pm)}(E)=\frac{\epsilon_{\pm}^{2}}{2\hbar\omega(1\mp\mathring{\kappa})^{2}},\quad\epsilon_{\pm}=\frac{E}{\hbar\omega}-1\pm\frac{1}{2}-\mathring{\kappa}$ (79)
-$\delta g_{L}^{(\pm)}(E)=\frac{2}{\hbar\omega(1\mp\mathring{\kappa})^{2}}\left(\epsilon_{\pm}\cot\frac{2\pi L}{1\mp\mathring{\kappa}}\sin\frac{2\pi L}{1\mp\mathring{\kappa}}\epsilon_{\pm}+\mathop{\rm cosec}\nolimits^{2}\frac{2\pi L}{1\mp\mathring{\kappa}}\cos\frac{2\pi L}{1\mp\mathring{\kappa}}\epsilon_{\pm}\right),$ (80)
-$\delta g_{N}^{(0)}(E)=-\frac{1}{\hbar\omega}\mathop{\rm cosec}\nolimits^{2}(\mathring{\kappa}\pi N)\cos(2\pi N\epsilon_{\pm}),$ (81)
-which is equivalent to that in [36].
-It appears to provide the trace formula (1) with contributions of POs with primitive periods $T_{\pm}=2\pi/\omega(1\mp\mathring{\kappa})$ and $T_{0}=2\pi/\omega$, but the irrational “Maslov indices” and the amplitude factors have yet to be explained by the classical dynamics.
-Appendix B Other minor orbits with spin precession
+$(\mathring{\kappa}=\kappa s/\omega,\quad n=0,1,2,\cdots,\quad\mathring{l}=0,1,2,\cdots)$
+where $n,\mathring{l},\mathring{\jmath}$ are radial, orbital angular momentum, and total angular momentum quantum numbers, respectively. Then, the level density is expressed as
+$g(E)=\sum_{n,\mathring{l},\mathring{\jmath}=\mathring{l}\pm 1/2}(2\mathring{\jmath}+1)\delta(E-E_{n\mathring{l}\mathring{\jmath}}).$ (73)
+By applying the Poisson’s summation formula to the above double sum, the level density is transformed into the form of summation over PO contributions as
+$g(E)=\bar{g}(E)+\sum_{\rm po}A_{\rm po}\cos\left(\tfrac{1}{\hbar}T_{\rm po}E-\tfrac{\pi}{2}\mu_{\rm po}\right).$ (74)
+The obtained formula includes contributions of POs with three primitive periods,
+$T_{0}=\frac{2\pi}{\omega},\quad T_{\pm}=\frac{2\pi}{\omega\mp\kappa s}.$
+In the direction from classical to quantum, application of the WKB method was considered in [20].
+By diagonalizing the Hamiltonians with respect to the spin channels, two adiabatic Hamiltonians $H_{\pm}=H_{0}\pm\kappa s|\bm{l}|$ are obtained.
+Circular orbits with periods $T_{\pm}$ are found in the Hamiltonian $H_{\pm}$, respectively, which are equivalent to the circular family with frozen spin discussed in Sec.
+III.
+Straight-line orbits with period $T_{0}$ are also found in both $H_{\pm}$ but they have $\bm{l}=0$ and lie on the mode conversion point $H_{+}=H_{-}$, hence the contribution to the level density cannot be evaluated.
+Apart from the mode conversion problem, they noticed another difficulty in classical-quantum correspondence.
+Degeneracy of the PO family is related to the semiclassical order of its contribution to the level density.
+In general, the amplitude factor $A_{\rm po}$ in the trace formula is proportional to $\hbar^{-K/2}$ for the $K$-parameter families.
+In the exact trace formula, the lowest order term in the amplitude $A_{\rm po}$ for PO with periods $T_{\pm}$ is proportional to 2superscriptPlanck-constant-over-$\hbar^{-2}$ suggesting $K=4$ families.
+However, the circular orbits in $H_{\pm}$ form only $K=2$ families and the semiclassical order of their contributions are inconsistent.
+Then, how about the Hamiltonian in the extended phase space taking the spin motion into account.
+The authors of Ref.
+[20] seem to have noticed about the existence of frozen-spin circular and diametric POs with period $T_{\pm}$ and $T_{0}$, respectively, but they didn’t discuss the matter further, probably because they do not have enough information on the classical POs.
+Now we have information on all the classical POs in the extended phase space.
+Unfortunately, the problem in semiclassical order of the contributing POs are not simply solved with the present result.
+The orbits having the periods $T_{\pm}$ are only the frozen-spin circular orbits that form $K=2$ families as well as in the WKB method.
+However, they encounter successive bifurcations with varying energy, which might enhance the contribution to the level density.
+Bridge orbits between circular POs have $K=3$ and their periods vary from $T_{+}$ and $T_{-}$ with increasing energy, but their contributions are not existing in the current form of the exact trace formula (74).
+There are two kinds of POs having the period $T_{0}$: $K=2$ family of diametric orbit with frozen spin, and $K=4$ family of ‘4D’ orbits which emerge from the second bifurcations of the bridge POs.
+In the exact formula, the amplitude of PO with period $T_{0}$ suggests a $K=2$ family, hence a kind of cancellation between the $K=4$ families seems to be necessary.
+It is an interesting open question if the above difficulties can be overcome and a more appropriate quantum-classical correspondence can be obtained by the use of trace formula in the extended phase space with suitable uniform approximations[27, 28], and also by another way of resummation of quantum numbers from the quantum spectrum (72) to the formula (74).
+Appendix Other minor orbits with spin precession
 For the sake of completeness, I consider here the $l_{z}=0$ POs with spin precession, which are not considered in the main part because of their insignificant contribution to the trace formula.
@@
 one has
-$U_{0}=0,\quad|Z_{+}|^{2}-|Z_{-}|^{2}=\frac{s_{\perp}}{M\omega},\quad|Z_{+}|^{2}+|Z_{-}|^{2}=\frac{E-\kappa s_{\perp}^{2}}{M\omega^{2}}.$ (82)
+$U_{0}=0,\quad|Z_{+}|^{2}-|Z_{-}|^{2}=\frac{s_{\perp}}{M\omega},$
+$|Z_{+}|^{2}+|Z_{-}|^{2}=\frac{E-\kappa s_{\perp}^{2}}{M\omega^{2}}.$ (75)
 For $Z_{-}=0$, the periodicity condition is
@@
 $T=\frac{2\pi k}{\omega}.$
-This condition is satisfied only by specific values of $\kappa$ and certain spin orientations for which $\kappa s_{z}$ and $\kappa s_{\perp}$ are both commensurable with $\omega$, but the orbit can exist for continuous values of energy with coefficients
-$|Z_{\pm}|^{2}=\frac{1}{2}\left(\frac{E-\kappa s_{\perp}^{2}}{M\omega^{2}}\pm\frac{s_{\perp}}{M\omega}\right),\quad E\geq E_{\rm bif}=\omega s_{\perp}+\kappa s_{\perp}^{2}.$
+This condition is satisfied only by specific values of $\kappa$ and certain spin orientations for which $\kappa s_{z}$ and $\kappa s_{\perp}$ are both commensurable with $\omega$, but once the orbit emerges, it can exist for any higher energy $E$ with the coefficients
+$|Z_{\pm}|^{2}=\frac{1}{2}\left(\frac{E-\kappa s_{\perp}^{2}}{M\omega^{2}}\pm\frac{s_{\perp}}{M\omega}\right),$
+$E\geq E_{\rm bif}=\omega s_{\perp}+\kappa s_{\perp}^{2}.$
 $k=m\omega/\kappa s_{z}>\omega/\kappa s$ is a large integer, which again gives a long PO.
@@
 In the case of $(n,m,k)=(25,1,24)$, for instance, one sees
-$s_{z}=s_{\perp}=\frac{s}{\sqrt{2}},\quad\kappa=\frac{\omega}{24s_{z}}\fallingdotseq 0.118\,\frac{\omega}{\hbar},\quad E_{\rm bif}\fallingdotseq 0.368\,\hbar\omega.$
+$s_{z}=s_{\perp}=\frac{s}{\sqrt{2}},$
+$\kappa=\frac{\omega}{24s_{z}}\fallingdotseq 0.118\,\frac{\omega}{\hbar},\quad E_{\rm bif}\fallingdotseq 0.368\,\hbar\omega.$